OpenStudy (anonymous):

If f(x) = x2 + x and g(x) = x - 3, find f(g(7)). 20 28 32 53

OpenStudy (anonymous):

\[g(7) = 7 - 3 =4\] \[f(g(7)) = f(4) = 4^{2} + 4 = 16 + 4 =20\] So the answer is "20".